# VCR: Visual Caption Restoration

### v1

@@ 1 Introduction
Recent advancements in large language models, such as ChatGPT [50, 49] and Llama [58], have spurred significant interest and progress in the field of vision-language models. With models like GPT-4V [49] and Llava [37, 38, 39] blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
Amidst this evolving landscape, we introduce Visual Caption Restoration (VCR), a novel vision-language task designed to challenge existing models uniquely. VCR challenges these models to restore obscured texts within images, a task that demands an intricate synthesis of text, vision, and text embedded in the image. The VCR task is grounded in two key insights: (1) text embedded within images, with its characteristics different from common visual elements, represents a distinct modality that requires careful alignment of vision, textual data, and the structure of written texts, and (2) neuroscience findings that suggest that humans are proficient in recognizing partially occluded objects through sophisticated visual and cognitive processes [57, 51, 59, 16, 34]. By leveraging these insights, VCR seeks to explore how well vision-language models can handle texts embedded within images, aligning visual elements and natural language to mimic human-like multi-modal understanding and recognition.
@@
To address these limitations, our VCR task builds on the premise that effective text restoration from images requires an integrated understanding beyond current VQA benchmarks’ capabilities. For instance, in extreme cases, models rely on existing Optical Character Recognition (OCR) system to extract text from documents [54, 7]. The extracted text is then used as context for generating answers, without a true semantic alignment between the text and the visual elements of the document. This approach, while effective in simple scenarios, falls short in more complex settings where text is intricately woven into the visual narrative of the image.
@@
By releasing VCR-wiki and the accompanying dataset construction code, we aim to stimulate further research in this area, encouraging the development of models that can more adeptly navigate the nuanced landscape of the restoration of text embedded in images. This effort aligns with the broader goal of advancing vision-language models to achieve deeper, more intuitive understanding of multimedia content, bridging the gap between human and machine perception. Code is available at https://github.com/tianyu-z/VCR.
@@
Conduct empirical evaluations showing significant gaps between current models and human performance on the VCR task. This highlights the effectiveness of VCR for assessing advancements in VLMs, and underscores the necessity for innovative model architectures and training techniques.
@@ 2 VCR Task Description
For better clarity, we define text embedded in image () as text incorporated within the image. The term visual image (VI) pertains to the portion of the image that excludes the text embedded in the image. The string text (ST) is not part of the image itself, but is associated with it as a distinct textual element. It is usually the question prompt in the form of natural language, e.g., ‘What are the covered texts in the image? Please only guess the covered texts without outputting an explanation.’. Consequently, an element of a Visual Caption Restoration (VCR) task can be expressed as , where is represented as string and both and are presented in image form. This notation does not imply that and can be physically separated into two distinct image components. Instead, this definition is adopted merely to facilitate a clearer explanation of the concepts involved. Please refer to Figure 2 for an illustration of , , and .
A1 Existing tasks that are similar to VCR are the tasks of VQA and OCR. VQA takes as input images and a natural language question, and generates a free-form response. As the ground-truth response is not unique, evaluating VQA poses a major challenge. In contrast to VQA, OCR is a task where the ground truth responses are unique: OCR takes as input complete characters in image form and outputs a string representing the characters in the image, without considering the image context. Models pretrained with OCR are able to retrieve texts embedded in the input image, even if they are incomplete or vague. However, as the vagueness or occlusion of the textual components of the image increases, retrieving the original text without considering the remaining non-textual image context becomes harder, and OCR is no longer a good approach. VCR bridges the gap between OCR and VQA: it reconstructs the unique text found in the image while also considering the visual context of the rest of the image.
Figure 2 is an example VCR task in the hard mode, and Figure 1 shows an example VCR task in the easy mode. Although humans can still fill the blanks easily in the hard mode, it is nearly impossible for models with only OCR capabilities to recover the covered texts without using the context. This is because the pixel-level hints of single characters no longer correspond to a unique solution.
@@
The first aspect of importance stems from neuroscience discoveries about human cognitive abilities to recognize partially occluded objects [16, 34]. Although existing models can recognize objects and texts in images, they often struggle with the complexity of occluded objects due to significant information loss in the images. In contrast, humans excel at filling in missing information using a combination of low-level visual processing and high-level cognitive functions, such as those managed by the prefrontal cortex. This cortical area is known to handle complex cognitive processes like decision-making and memory retention, essential for integrating fragmented visual inputs into coherent objects. We believe that the occlusion restoration task serves as a probe that can effectively distinguish a model’s low-level recognition and high-level cognition involving reasoning. Besides, understanding these neural mechanisms can inspire new algorithms capable of mimicking human-like perception and interpretation in dynamic, real-world conditions where occlusion is common.
The second aspect underscores the unique challenge presented by the VCR task, distinguishing it significantly from existing benchmarks like traditional VQA or occluded object restoration task. By occluding texts instead of common visual objects, VCR targets the models’ text-image alignment capability, which is one of the major challenges for vision-language models. VCR mandates that models align textual and visual information in a manner that replicates human-like understanding involving the utilization of both textual and visual clues. This task necessitates a deep integration of visual (), embedded textual (), and contextual interpretation across modalities, pushing beyond simple text extraction as performed in OCR tasks. In OCR, the focus is primarily on recognizing visible characters, often without a need to understand their contextual relevance within the image narrative. In contrast, VCR introduces complexity by requiring the model to use available partial texts and the visual context collaboratively to reconstruct the obscured content accurately. This not only tests the model’s ability to process - modalities effectively, but also challenges it to maintain internal consistency, akin to human cognitive processes where context and visual clues guide understanding and response. Besides, the difficulty of the task can be adjusted by altering the extent of text occlusion, offering a scalable and flexible framework for systematically enhancing model capabilities in text-visual alignment and semantic comprehension. This rigorous testing ground will help evolve vision-language models to better grasp the nuanced interplay between text and imagery.
@@ 3 Dataset Creation
The VCR task requires aligning visual images () with text embedded in images (). Therefore, the dataset creation process relies on a set of highly correlated image-text pairs. We utilize the primary images and their corresponding captions from Wikipedia as the data source111Datasource: https://huggingface.co/datasets/wikimedia/wit_base. to create VCR-wiki, a Wikipedia-based VCR dataset. The creation pipeline for VCR-wiki is shown in Figure 2. Before constructing the dataset, we first filter out the instances with sensitive content, including NSFW and crime-related terms, to mitigate AI risk and biases.
The VCR-wiki dataset is formatted as a VQA task, where each instance includes an image, a question, and a ground-truth answer. The images are synthesized from the text-image pairs by stacking the image () with its corresponding text description () vertically, mimicking the format of a captioned image. This stacked image is referred to as a stacked + image. Each + image is resized to a width of 300 pixels. To prevent excessive image height, we truncate to a maximum of five lines. We filter the dataset to exclude instances with + images exceeding 900 pixels in height to avoid drastic resolution change during data preprocessing.
Within , we use spaCy to randomly select several 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much professional domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates and time labeled by spaCy, and the total masked token does not exceed 50% of the tokens in the caption. We exclude all instances that do not have a single eligible 5-gram. The selected 5-grams are partially obscured by a white rectangle that reveals only the top and bottom parts of the text, with the proportion of coverage varying to adjust task difficulty. Furthermore, to assess the impact of on model performance, we create an ablation for each image, maintaining the resolution of the + image, but retaining only the part in the center of the image.
The VCR task involves a predefined question that prompts the model to produce the text obscured in the image. The ground-truth answer corresponds to the caption displayed in the uncovered portion of the stacked image. Due to the extensive availability of vision-language models and a significant user base in both English and Chinese, we have chosen to develop the dataset in these two languages. For each language, we meticulously select the covering proportion to create two task variants: (1) an easy version, where the task is easy for native speakers but open-source OCR models almost always fail, and (2) a hard version, where the revealed part consists of only one to two pixels for the majority of letters or characters, yet the restoration task remains feasible for native speakers of the language.
@@ 3.1 Dataset Format and Statistics
The VCR dataset comprises four configurations: English Easy, English Hard, Chinese Easy, and Chinese Hard. Each configuration can be further divided into training, validation, and test splits. Validation and test splits contain 5,000 entities each. The training set for English configurations and the Chinese configurations contain and instances, respectively, which can be used for model continuous pretraining. We include more detailed statistics of the dataset in Table 1.
@@ 4.1 Models
We evaluate open-source models with the best performance on the OpenVLM Leaderboard222We selected the highest-performing open-source models with fewer than 40 billion parameters from the OpenVLM Leaderboard as of May 23, 2024. Details are available at: https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. and state-of-the-art Chinese VLM models. The evaluated models include InternVL-Chat-V1.5[11], MiniCPM-Llama3-V2.5 [25], InternLM-XComposer2-VL-7B [14], CogVLM2-Llama3-19B-Chat [63], Idefics2-8B [32], Yi-VL-34B [1], Yi-VL-6B [1], Qwen-VL-Chat [4], DeepSeek-VL-7B-Chat [42], DeepSeek-VL-1.3B-Chat [42], TextMonkey [40] and DocOwl-1.5 [22]. Out of these models, Idefics2-8B is an English-only model, and CogVLM2-Llama3-19B-Chat has its Chinese variant, CogVLM2-Llama3-19B-Chinese-Chat. Please refer to Table 2 for the model specifications.
@@
More specifically, We fine-tune CogVLM2-Llama3-19B-Chat and MiniCPM-Llama3-V2.5 on the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat and MiniCPM-Llama3-V2.5 on the Chinese Hard configuration. The models are finetuned using LoRA [23] with and . We adopt the schedule-free AdamW optimizer [12] with a learning rate . The effective batch size is 64. Each model is trained on the first 16000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
@@ 4.3 Experimental Results
Although achieving state-of-the-art performance on the Open VLM leaderboard, almost all the tested models achieve a low exact match accuracy on the English Easy configuration and fail on the other settings. The best open-source model across the 4 configurations (English Easy, English Hard, Chinese Easy and Chinese Hard) is CogVLM2-Llama3-Chat. This might be attributed to its pretraining process and the special architecture. We also notice that VI has negative impact to most models on the exact match scores () which means the image information is not properly utilized. The best performed open-source model, CogVLM2-Llama3-Chat and its finetuned version has positive except the Chinese Hard configuration. This indicates that information from VI could help improve the model performance on VCR.
For different languages, we noticed a large performance drop when testing the model in Chinese configurations, despite the fact that all the models claim to have basic English-Chinese duolingual capabilities. This is somehow surprising, since Chinese characters, due to their logographic nature, may exhibit a higher degree of recognizability compared to languages that utilize alphabetic scripts in one order [65, 73].
@|c|c|lC1.5cmcccccc|@
Lang
@@
EN Easy Closed-source models
@@
ZH Easy Closed-source models
@@
Moreover, we found that models, such as internlm-xcomposer2, are good at OCR and understanding image documents (as demonstrated by DocOwl 1.5 and TextMonkey) still have the potential to be improved in the VCR task. This highlights the unique and indispensable role of VCR in the current suite of benchmarks. Excelling in other document-related benchmarks does not guarantee similar performance in VCR tasks, emphasizing VCR’s distinct challenges and value.
@@
We evaluate every closed-source model with the first 500 examples in the test set. In English tasks, GPT-4o achieves the best scores among the models which have not been finetuned. Even as GPT-4 series and Claude support Chinese, we found that they cannot completely recognize Chinese words embedded in the image even without any occlusion. Thus, we do not test GPT-4 and Claude model series on Chinese tasks.
In the English configurations, closed-source models outperform all open-source models except CogVLM2, which indicates that model scaling might help improve performance on the VCR task. However, compared with the human evaluation results in Section 4.4, we notice a large performance gap, especially in the English Hard configuration. This shows significant room for improvement in the current state-of-the-art models.
Refer to Table A to compare open-source and closed-source models using the same 500 test cases.

### v2

@@ 1 Introduction
Recent advances in large language models, such as ChatGPT [51, 50] and Llama [62], have spurred significant interest and progress in the field of vision-language models. With models like GPT-4V [50] and Llava [38, 39, 40] blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
Amidst this evolving landscape, we introduce Visual Caption Restoration (VCR), a novel vision-language task designed to challenge existing models uniquely. VCR challenges these models to restore obscured texts within images, a task that demands an intricate synthesis of text, vision, and text embedded in the image. The VCR task is grounded in two key insights: (1) text embedded within images, with its characteristics different from common visual elements, represents a distinct modality that requires careful alignment of vision, textual data, and the structure of written texts, and (2) neuroscience findings that suggest that humans are proficient in recognizing partially occluded objects through sophisticated visual and cognitive processes [61, 52, 63, 16, 34]. By leveraging these insights, VCR seeks to explore how well vision-language models can handle texts embedded within images, aligning visual elements and natural language to mimic human-like multimodal understanding and recognition.
@@
To address these limitations, our VCR task builds on the premise that effective text restoration from images requires an integrated understanding beyond the capabilities of current VQA benchmarks. For example, in extreme cases, models rely on existing Optical Character Recognition (OCR) system to extract text from documents [56, 7]. The extracted text is then used as context for generating answers, without a true semantic alignment between the text and the visual elements of the document. This approach, while effective in simple scenarios, falls short in more complex settings where text is intricately woven into the visual narrative of the image.
@@
By releasing VCR-wiki and the accompanying dataset construction code, we aim to stimulate further research in this area, encouraging the development of models that can more adeptly navigate the nuanced landscape of the restoration of text embedded in images. This effort aligns with the broader goal of advancing vision-language models to achieve a deeper, more intuitive understanding of multimedia content, bridging the gap between human and machine perception. The code is available at https://github.com/tianyu-z/VCR.
@@
Conduct empirical evaluations that show significant gaps between current models and human performance on the VCR task. This highlights the effectiveness of VCR for assessing advancements in VLMs, and underscores the necessity for innovative model architectures and training techniques.
@@ 2 VCR Task Description
For better clarity, we define text embedded in image () as text incorporated within the image. The term visual image (VI) pertains to the portion of the image that excludes the text embedded in the image. The string text (ST) is not part of the image itself, but is associated with it as a distinct textual element. It is usually the question prompt in the form of natural language, for example, ‘What are the covered texts in the image? Please only guess the covered texts without outputting an explanation.’. Consequently, an element of a Visual Caption Restoration (VCR) task can be expressed as , where is represented as a string and both and are presented in image form. This notation does not imply that and can be physically separated into two distinct image components. Instead, this definition is adopted merely to facilitate a clearer explanation of the concepts involved. Please refer to Figure 2 for an illustration of , , and .
A1 Existing tasks that are similar to VCR are the tasks of VQA and OCR. VQA takes as input images and a natural language question and generates a free-form response. As the ground-truth response is not unique, evaluating VQA poses a major challenge. In contrast to VQA, OCR is a task where the ground-truth responses are unique: OCR takes as input complete characters in image form and outputs a string representing the characters in the image, without considering the image context. Models pretrained with OCR are able to retrieve texts embedded in the input image, even if they are incomplete or vague. However, as the vagueness or occlusion of the textual components of the image increases, retrieving the original text without considering the remaining nontextual image context becomes harder, and OCR is no longer a good approach. VCR bridges the gap between OCR and VQA: it reconstructs the unique text found in the image while also considering the visual context of the rest of the image.
Figure 2 is an example VCR task in hard mode, and Figure 1 shows an example VCR task in the easy mode. Although humans can still fill the blanks easily in the hard mode, it is nearly impossible for models with only OCR capabilities to recover the covered texts without using the context. This is because the pixel-level hints of single characters no longer correspond to a unique solution.
@@
The first aspect of importance stems from discoveries in neuroscience about human cognitive abilities to recognize partially occluded objects [16, 34]. Although existing models can recognize objects and texts in images, they often struggle with the complexity of occluded objects due to significant information loss in the images. In contrast, humans excel at filling in missing information using a combination of low-level visual processing and high-level cognitive functions, such as those managed by the prefrontal cortex. This cortical area is known to handle complex cognitive processes such as decision making and memory retention, which are essential for integrating fragmented visual input into coherent objects. We believe that the occlusion restoration task serves as a probe that can effectively distinguish low-level recognition and high-level cognition involving reasoning. In addition, understanding these neural mechanisms can inspire new algorithms capable of mimicking human-like perception and interpretation in dynamic, real-world conditions where occlusion is common.
The second aspect underscores the unique challenge presented by the VCR task, distinguishing it significantly from existing benchmarks, such as traditional VQA or the occluded object restoration task. By occluding texts instead of common visual objects, VCR targets the models’ text-image alignment capability, which is one of the major challenges for vision-language models. VCR mandates that models align textual and visual information in a manner that replicates human-like understanding involving the utilization of both textual and visual clues. This task requires a deep integration of visual (), embedded textual (), and contextual interpretation across modalities, pushing beyond simple text extraction as performed in OCR tasks. In OCR, the focus is primarily on recognizing visible characters, often without the need to understand their contextual relevance within the image narrative. In contrast, VCR introduces complexity by requiring the model to use available partial texts and the visual context collaboratively to reconstruct the obscured content accurately. This not only tests the model’s ability to process - modalities effectively, but also challenges it to maintain internal consistency, akin to human cognitive processes where context and visual clues guide understanding and response. Besides, the difficulty of the task can be adjusted by altering the extent of text occlusion, offering a scalable and flexible framework for systematically enhancing model capabilities in text-visual alignment and semantic comprehension. This rigorous testing ground will help evolve vision-language models to better grasp the nuanced interplay between text and imagery.
@@ 3 Dataset Creation
The VCR task requires aligning visual images () with text embedded in images (). Therefore, the dataset creation process relies on a set of highly correlated image-text pairs. We utilize the primary images and their corresponding captions from Wikipedia as the data source111Datasource: https://huggingface.co/datasets/wikimedia/wit_base. to create VCR-wiki, a Wikipedia-based VCR dataset. The pipeline for creating VCR-wiki is shown in Figure 2. Before constructing the dataset, we first filter out instances with sensitive content, including NSFW and crime-related terms, to mitigate AI risk and biases.
The VCR-wiki dataset is formatted as a VQA task, where each instance includes an image, a question, and a ground-truth answer. The images are synthesized from text-image pairs by stacking the image () with its corresponding text description () vertically, mimicking the format of a captioned image. This stacked image is referred to as a stacked + image. Each + image is resized to a width of 300 pixels. To avoid excessive image height, we truncate to a maximum of five lines. We filter the dataset to exclude instances with + images exceeding 900 pixels in height to avoid drastic resolution changes during data pre-processing.
Within , we use spaCy to randomly select several 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much professional domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates and time labeled by spaCy, and the total masked token does not exceed 50% of the tokens in the caption. We exclude all instances that do not have a single eligible 5-gram. The selected 5-grams are partially obscured by a white rectangle that reveals only the upper and lower parts of the text, with the proportion of coverage varying to adjust task difficulty. Furthermore, to assess the impact of on model performance, we create an ablation for each image, maintaining the resolution of the + image, but retaining only the part in the center of the image.
The VCR task involves a predefined question that prompts the model to produce the obscured text in the image. The ground truth answer corresponds to the caption displayed in the uncovered portion of the stacked image. Due to the extensive availability of vision-language models and a significant user base in both English and Chinese, we have chosen to develop the dataset in these two languages. For each language, we meticulously select the covering proportion to create two task variants: (1) an easy version, where the task is easy for native speakers but open-source OCR models almost always fail, and (2) a hard version, where the revealed part consists of only one to two pixels for the majority of letters or characters, yet the restoration task remains feasible for native speakers of the language.
@@ 3.1 Dataset Format and Statistics
The VCR dataset comprises four configurations: English Easy, English Hard, Chinese Easy and Chinese Hard. Each configuration can be further divided into training, validation, and test splits. The validation and test splits contain 5,000 entities each. The training set for English configurations and Chinese configurations contains and instances, respectively, which can be used for model continuous pretraining. We include more detailed statistics of the dataset in Table 1.
@@ 4.1 Models
We evaluate several most advanced proprietary models with their provided APIs. The evaluated models include GPT-4o (gpt-4o-2024-0513), GPT-4 Turbo (gpt-4-turbo-2024-04-09), GPT-4V (gpt-4-1106-vision-preview) [51, 50], Claude 3 Opus (claude-3-opus-20240229), Claude 3.5 Sonnet (claude-3-5-sonnet-20240620) [2], Gemini 1.5 pro (gemini-1.5-pro-001) [59], Reka Core (reka-core-20240501) [60], and Qwen-VL-Max (tested on May 2024) [4].
@@
We evaluate open-source models with the best performance on the OpenVLM Leaderboard222We selected the highest-performing open-source models with fewer than 40 billion parameters from the OpenVLM Leaderboard as of May 23, 2024. Details are available at https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. and state-of-the-art Chinese VLM models. The evaluated models include InternVL-Chat-V1.5[11], MiniCPM-Llama3-V2.5 [25], InternLM-XComposer2-VL-7B [14], CogVLM2-Llama3-19B-Chat [67], Idefics2-8B [32], Yi-VL-34B [1], Yi-VL-6B [1], Qwen-VL-Chat [4], DeepSeek-VL-7B-Chat [43], DeepSeek-VL-1.3B-Chat [43], Monkey [41, 35] and DocOwl-1.5 [22]. Out of these models, Idefics2-8B is an English-only model, and CogVLM2-Llama3-19B-Chat has its Chinese variant, CogVLM2-Llama3-19B-Chinese-Chat. Please refer to Table 2 for the model specifications.
@@
More specifically, we fine-tune CogVLM2-Llama3-19B-Chat and MiniCPM-Llama3-V2.5 in the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat and MiniCPM-Llama3-V2.5 on the Chinese Hard configuration. The models are finetuned using LoRA [23] with and . We adopt the schedule-free AdamW optimizer [12] with a learning rate . The effective batch size is 64. Each model is trained on the first 16,000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
@@ 4.3 Experimental Results
Although achieving state-of-the-art performance on the Open VLM leaderboard, almost all the tested models achieve a low exact match accuracy in the English Easy configuration and fail on the other settings. The best open-source model across the 4 configurations (English Easy, English Hard, Chinese Easy, and Chinese Hard) is CogVLM2-Llama3-Chat. This might be attributed to its pretraining process and the special architecture. We also notice that VI has a negative impact for most models on the exact match scores (), which means that the image information is not properly utilized. The best performed open-source model, CogVLM2-Llama3-Chat, and its fine-tuned version have positive , except for the Chinese Hard configuration. This indicates that information from VI could help improve the model performance on VCR.
For different languages, we noticed a large performance drop when testing the model in Chinese configurations, despite the fact that all models claim to have basic English-Chinese duolingual capabilities. This is somehow surprising, since Chinese characters, due to their logographic nature, may exhibit a higher degree of recognizability compared to languages that use alphabetic scripts in one order [69, 79].
@|c|c|p4cmC2cmC2.3cmC2.3cmC2cmC2.3cmC2.3cmC2cm|@
Language
@@
English Easy Closed-source models
@@
Chinese Easy Closed-source models
Claude 3 Opus - -0.1 1.49
Claude 3.5 Sonnet - 0.2 0.03
Gemini 1.5 Pro - 0.6 -0.37
GPT-4o - -7.58 -9.19
GPT-4 Turbo - 0.1 1.45
@@
Claude 3 Opus - 0.2 1.13
Claude 3.5 Sonnet - 0.2 1.63
Gemini 1.5 Pro - 0.2 0.07
GPT-4o - 0.4 -0.17
GPT-4 Turbo - -0.2 1.72
@@
Moreover, we found that models, such as internlm-xcomposer2, are good at OCR and understanding image documents (as demonstrated by DocOwl 1.5 and Monkey) still have the potential to be improved in the VCR task. This highlights the unique and indispensable role of VCR in the current suite of benchmarks. Excelling in other document-related benchmarks does not guarantee similar performance in VCR tasks, emphasizing VCR’s distinct challenges and value.
@@
We evaluate every closed-source model with the first 500 examples in the test set. In English tasks, GPT-4o scores the best among the models that have not been finetuned. Even though GPT-4 series support Chinese, we found that GPT-4V (gpt-4-1106-vision-preview) is not able to recognize Chinese characters embedded in the image even without any occlusion. Thus, we do not test GPT-4V on Chinese tasks.
In English configurations, closed-source models outperform all open-source models except CogVLM2, which indicates that model scaling might help improve performance on the VCR task. However, compared with the human evaluation results in Section 4.4, we notice a large performance gap, especially in the English Hard configuration. This shows significant room for improvement in the current state-of-the-art models.
Refer to Table A to compare open and closed source models using the same 500 test cases.